Where in the world do I start?
On June 1, 1998, the average age of the 33 employees at Acme Cement was 47 years. On June 1, 1999, three of the staff aged 65,58,and 62 retired and were replaced by four employees aged 24,31,26 and 28. What was the average age of the employees at Acme Cement on June 1, 1999?

Answer by checkley71(8403)

TOTAL AGE OF 33 EMPLOYEES IS 33*47=1551 SO WE SUBTRACT 185 YEARS FOR THE 3 RETIREES AND ADD 109 FOR THE 4 NEW EMPLOYEES WHICH GIVES US A NEW TOTAL OF
1551-185+109=1475 NOW WE NEED TO DIVIDE THIS NEW TOTAL BY 34 EMPLOYEES OR
1475/34=43.38 AVERAGE AGE.